OpenStudy (vivek3461):

Kindly Clarify this... when i try to find the solution of a first order autonomous differential equation by using the integrating factor method used for linear first order DEs i m getting only a constant solution. What about the other general solutions??

OpenStudy (zehanz):

An example of a forst order autonomous d.e. is: y'= 2y+3, or \(\frac{dy}{dx}=2y+3\), so assuming y is a function of x. The integrating factor is \(e^{\int2dx}=e^{2x}\). General solution: \(y=e^{2x}(\int e^{-2x}\cdot3dx+C)\), where C is a real constant. This can be simplified to: \(y=e^{2x}(-\frac{3}{2}e^{-2x}+C)=-\frac{3}{2}+Ce^{2x}\). The constant solution occurs for C=0.

OpenStudy (dumbcow):

holy sh*t, this is a year old

OpenStudy (zehanz):

Thanks anyway :)

OpenStudy (vivek3461):

Thank you ZeHanz....

OpenStudy (zehanz):

yw!

OpenStudy (zehanz):

Hope my answer wasn't too late :)

OpenStudy (vivek3461):

Late is always better than never...:-)